OpenStudy (anonymous):

The formula to convert Fahrenheit to Celsius is C=5/9(F=32). Convert 12C to Fahrenheit. Round to the nearest degree.

OpenStudy (anonymous):

Pretty simple, plug in 12 for C and Solve for F.

OpenStudy (anonymous):

what about 5/9

OpenStudy (anonymous):

\[F=\frac{ 9 }{ 5 }C+32\]

OpenStudy (anonymous):

i dont understand???

OpenStudy (anonymous):

\[C=\frac{ 5 }{ 9 }(F-32)\] rearrange the equation so you F is alone on 1 side. \[9C=5(F-32)\] \[\frac{ 9 }{ 5 }C=F-32\] \[\frac{ 9 }{ 5 }C+32=F\] makes sense now? and now to get 12C into Farenheit, simple make 12=C and solve for F.

OpenStudy (anonymous):

so the answer would be 54F

OpenStudy (anonymous):

ot 79F

OpenStudy (anonymous):

or*

OpenStudy (anonymous):

53.6=F